given that v²=x²+y²+z² prove that d²v/dx² +d²v/dy² +d²v/dz² =2/v
1
Expert's answer
2020-04-13T19:12:53-0400

"v^2=x^2+y^2+z^2"

"v" is symmetrical for "x,y,z" . Let's found "\partial v/\partial x" :

"v=\sqrt{x^2+y^2+z^2}"

"{\frac {\partial v} {\partial x}}={\frac {2x} {2\sqrt{x^2+y^2+z^2}}}={\frac {x} {v}}"

Then "{\frac {\partial v} {\partial y}}={\frac {y} {v}}" and "{\frac {\partial v} {\partial z}}={\frac {z} {v}}"

Let's found second order derivatives

"{\frac {\partial^2 v} {\partial x^2}}={\frac {v-x{\frac {\partial v} {\partial x}}} {v^2}}={\frac {v-{\frac {x^2} {v}}} {v^2}}={\frac {v^2-x^2} {v^3}}"

Then "{\frac {\partial^2 v} {\partial y^2}}={\frac {v^2-y^2} {v^3}}" and "{\frac {\partial^2 v} {\partial z^2}}={\frac {v^2-z^2} {v^3}}"

Let's prove our statement

"{\frac {\partial^2 v} {\partial x^2}}+{\frac {\partial^2 v} {\partial y^2}}+{\frac {\partial^2 v} {\partial z^2}}={\frac {v^2-x^2} {v^3}}+{\frac {v^2-y^2} {v^3}}+{\frac {v^2-z^2} {v^3}}={\frac {3v^2-x^2-y^2-z^2} {v^3}}={\frac {2v^2} {v^3}}={\frac {2} {v}}"

Proved